A stone dropped from the top of a building covers 36percent of the total distance in the last sec of it`s motion .the height of the building is?

Grade:12

1 Answers

Arun
6 years ago
Dear Radha
the time taken for a freely falling body to cover n %  in the last second is ?n...
That is time = under root 'n' 
Where n = percent of dust. covered in last second....
So time = 6sec
Now....S = ut + 1/2 at.... 
But here.  u = 0.. So.... S= 1/2 g 6?2.    
If g = 10 then dust. Or height = 5 × 36 = 180 meters...
 
Regards
Arun (askIITians forum expert)